# Newton’s method for nonlinear mappings into vector bundles

### v1

@@ 5.1. Local norms and a metric on manifolds
Let $x_{\star}\in\mathcal{X}$ and let $R$ be a $C^{1}$-retraction. Then there exists a neighborhood $V\subseteq T\mathcal{X}$ of $(x_{\star},0)\in T\mathcal{X}$ such that $R:V\rightarrow R(V)$ is a diffeomorphism.
Finally, we can relate retractions, local norms and the metric on $\mathcal{X}$ defined by the local norms.
@@
To $x_{\star}\in\mathcal{X}$ and $01>\varepsilon>$ 0 there exists a neighborhood $V$ of $x_{\star}$ such that for all $x,y,z\in V$ the following holds:
@@
Choose a chart $(U,\phi)$ of $\mathcal{X}$ at $x_{\star}$ and denote by $\Phi:\pi^{-1}(U)\to U\times\mathbb{X}$ the natural vector bundle chart on the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$. Let $x_{\star,\phi},\ x_{\phi},\ y_{\phi}$ and $z_{\phi}$ be representatives of $x_{\star},x,y$ and $z$ in the chart. Let $R_{\phi}^{-1}$ be the chart representation of $R^{-1}$. We show that
$\lim_{x_{\phi},y_{\phi},z_{\phi}\rightarrow x_{\star,\phi}}\frac{\|R_{\phi}^{-1}(z_{\phi},x_{\phi})-R_{\phi}^{-1}(z_{\phi},y_{\phi})\|_{\Phi,x_{\star}}}{\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}}}=1.$
Since $R^{-1}$ is continuously differentiable, $R^{-1}$ is also strictly differentiable. Thus, by additionally using the properties of a retraction, there exists a neighborhood $V$, such that for arbitrary $\varepsilon>0$ we obtain, denoting by $r_{R^{-1}}(x_{\phi},y_{\phi})$ the remainder term of strict differentiability:
$\|R_{\phi}^{-1}(z_{\phi},x_{\phi})-R_{\phi}^{-1}(z_{\phi},y_{\phi})\|_{\mathbb{X}}$ $=\|(R_{\phi}^{-1})^{\prime}(x_{\star,\phi},x_{\star,\phi})(x_{\phi}-y_{\phi})+r_{R^{-1}}(x_{\phi},y_{\phi})\|_{\mathbb{X}}$
$=\|R_{\phi}^{\prime}(R_{\phi}^{-1}(x_{\star,\phi},x_{\star,\phi}))(x_{\phi}-y_{\phi})+r_{R^{-1}}(x_{\phi},y_{\phi})\|_{\mathbb{X}}$
$=\|x_{\phi}-y_{\phi}+r_{R^{-1}}(x_{\phi},y_{\phi})\|_{\mathbb{X}}$
$\leq\|x_{\phi}-y_{\phi}\|_{\mathbb{X}}+\varepsilon\|x_{\phi}-y_{\phi}\|_{\mathbb{X}}.$
Using that $\|\cdot\|_{\Phi,x_{\star}}$ is equivalent to $\|\cdot\|_{\mathbb{X}}$ and by possibly shrinking $V$, we obtain the desired assertion by Prop. 5.5. ∎
@@ 5.2. Newton- and Strict Differentiability in Geometric Form
In order to analyse local convergence of Newton’s method, we want to prove a result about a Newton- or strictly differentiable mapping $F:\mathcal{X}\to\mathcal{E}$ between a manifold $\mathcal{X}$ and a vector bundle $p:\mathcal{E}\to\mathcal{M}$, which can be stated independently of the choice of charts.
@@
$y=p\circ F:\mathcal{X}\to\mathcal{M},\;y(x):=p(F(x))\in\mathcal{M},$
which computes the base point of $F(x)\in\mathcal{E}$ for any $x\in\mathcal{X}$, is locally Lipschitz continuous. We may use the linear map $V^{-1}_{y}(y_{0})\in L(E_{y_{0}},E_{y})$ given by a vector back-transport $V_{y}^{-1}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ to transport elements $F(x_{0})\in E_{y_{0}}$ into the fibre $E_{y}$.
@@
Let $F:\mathcal{X}\to\mathcal{E}$. Let $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ be a vector back-transport, $Q:T\mathcal{E}\to\mathcal{E}$ be a connection map and $R:T\mathcal{X}\to\mathcal{X}$ be a retraction.
@@
    Let $F:\mathcal{X}\to\mathcal{E}$ be Newton-differentiable at a point $\xi\in\mathcal{X}$ with respect to a fibrewise linear mapping $F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{E}))$. If $V^{-1}_{y}$ is consistent with $Q$ or $F(\xi)=0$, then it holds
    $\underset{x\to\xi}{\lim}\frac{\|(Q_{F(x)}\circ F^{\prime}(x))(R_{x}^{-1}(x)-R_{x}^{-1}(\xi))-(F(x)-V_{y(x)}^{-1}(y(\xi))F(\xi))\|_{E_{y(x)}}}{d_{\mathcal{X}}(x,\xi)}=0.$
@@
    Let $F$ be strictly differentiable at $a\in\mathcal{X}$. If $V^{-1}_{y}$ is consistent with $Q$ or $F(a)=0$ then for every $\varepsilon>0$ there exists a neighborhood $W$ of $a$ such that for all $x,\xi\in W$ it holds
    $\|(Q_{F(a)}\circ F^{\prime}(a))(R_{a}^{-1}(x)$ $-R_{a}^{-1}(\xi))-(V^{-1}_{y(a)}(y(x))F(x)-V_{y(a)}^{-1}(y(\xi))F(\xi))\|_{E_{y(a)}}$
    $<\varepsilon\|R_{a}^{-1}(x)-R_{a}^{-1}(\xi)\|_{a}.$
@@
Choose a chart $\phi$ of $\mathcal{X}$ and a vector bundle chart $\eta$ of the vector bundle $\mathcal{E}$. In the following, we use $\phi$ or $\eta$ as an index to denote the chart representation of elements of $\mathcal{X}$ or $\mathcal{E}$, respectively. We begin by discussing the numerators in $(i)$ and $(ii)$ in a unified way. Then for the proof of $(i)$ we will fix $\xi$ and consider the case $x=a$ and $x\to\xi$. For the proof of $(ii)$ we will fix $a$ and consider $x,\xi\to a$.
Let $R_{a,\phi}^{-1}$ be a representation of the inverse retraction in charts. Since $R_{a}^{-1}$ is a $C^{1}$-mapping and using $(R_{a}^{-1})^{\prime}(a)=Id_{T_{a}\mathcal{X}}$, we obtain
$R_{a,\phi}^{-1}(x_{\phi})-R_{a,\phi}^{-1}(\xi_{\phi})=(R^{-1}_{a,\phi})^{\prime}(a_{\phi})(x_{\phi}-\xi_{\phi})+r_{R}=(x_{\phi}-\xi_{\phi})+r_{R}$
where $r_{R}$ is a remainder term. Setting $\delta x_{\phi}=x_{\phi}-\xi_{\phi}$ and $\delta y_{\eta}:=y(x)_{\eta}-y(\xi)_{\eta}\in\mathbb{M}$, we obtain the chart representation of
$(Q_{F(a)}\circ F^{\prime}(a))(R_{a}^{-1}(x)-R_{a}^{-1}(\xi))$
@@
$(F_{\eta})^{\prime}(a_{\phi})(\delta x_{\phi}+r_{R})-B_{y(a),\eta}(F_{\eta}(a_{\phi}))\delta y_{\eta}.$ (11)
Next, we consider the representation of
$V_{y(a)}^{-1}(y(x))F(x)-V_{y(a)}^{-1}(y(\xi))F(\xi),$
which is of the form
$V^{-1}_{y(a),\eta}(y(x)_{\eta})F_{\eta}(x_{\phi})-V^{-1}_{y(a),\eta}(y(\xi)_{\eta})F_{\eta}(\xi_{\phi}).$ (12)
Using the differentiability of the mapping $V_{y,\eta}^{-1}:\mathbb{M}\to L(\mathbb{E},\mathbb{E})$, we can write:
$V_{y(a),\eta}^{-1}(y(x)_{\eta})-V_{y(a),\eta}^{-1}(y(\xi)_{\eta})=(V_{y(a),\eta}^{-1})^{\prime}(y(a)_{\eta})\delta y_{\eta}+r_{V}.$
Inserting this into (12) we obtain by linearity of $V_{y}^{-1}(y)$
$((V_{y(a),\eta}^{-1})^{\prime}(y(a)_{\eta})\delta y_{\eta})F_{\eta}(\xi_{\phi})+V_{y(a),\eta}^{-1}(y(x)_{\eta})(F_{\eta}(x_{\phi})-F_{\eta}(\xi_{\phi}))+r_{V}F_{\eta}(\xi_{\phi}).$ (13)
Writing the difference $F_{\eta}(x_{\phi})-F_{\eta}(\xi_{\phi})$ as
$F_{\eta}(x_{\phi})-F_{\eta}(\xi_{\phi})=(F_{\eta})^{\prime}(a_{\phi})\delta x_{\phi}+r_{F}.$
@@
$(Id_{E_{y(a)}}-V_{y(a),\eta}^{-1}(y(x)_{\eta}))(F_{\eta})^{\prime}(a_{\phi})\delta x_{\phi}$
$-B_{y(a),\eta}(F_{\eta}(a_{\phi}))\delta y_{\eta}-((V_{y(a),\eta}^{-1})^{\prime}(y(a)_{\eta})\delta y_{\eta})F_{\eta}(\xi_{\phi})$
$+(F_{\eta})^{\prime}(a_{\phi})r_{R}-V_{y(a),\eta}^{-1}(y(x)_{\eta})r_{F}-r_{V}F_{\eta}(\xi_{\phi}).$
Now consider the different cases.
$(i)$: Set $a=x$ and let $F$ be Newton-differentiable at $\xi$. Since $V_{y}^{-1}(y)=Id_{E_{y}}$ for $y\in\mathcal{M}$ and $x=a$ our representation reduces to
$-B_{y(x),\eta}(F_{\eta}(x_{\phi}))\delta y_{\eta}-((V_{y(x),\eta}^{-1})^{\prime}(y(x)_{\eta})\delta y_{\eta})F_{\eta}(\xi_{\phi})+(F_{\eta})^{\prime}(x_{\phi})r_{R}-r_{F}-r_{V}F_{\eta}(\xi_{\phi}).$
Since $R_{x}^{-1}$ is continuously differentiable, it is also Newton-differentiable and thus the remainder term $r_{R}$ is of order $o(\|x_{\phi}-\xi_{\phi}\|_{\mathbb{X}})$ as $x_{\phi}\to\xi_{\phi}$. Since $F$ is Newton-differentiable with respect to $F^{\prime}$, $(F_{\eta})^{\prime}(x_{\phi})$ is locally bounded. Thus, the remainder term $(F_{\eta})^{\prime}(x_{\phi})r_{R}$ is still of order $o(\|x_{\phi}-\xi_{\phi}\|_{\mathbb{X}})$ as $x_{\phi}\to\xi_{\phi}$. Since $F$ is Newton-differentiable $r_{F}$ is also of order $o(\|x_{\phi}-\xi_{\phi}\|_{\mathbb{X}})$ as $x_{\phi}\to\xi_{\phi}$. The remainder term $r_{V}$ depends on $y(x)_{\eta}$ and $y(\xi)_{\eta}$ and is of order $o(\|y(x)_{\eta}-y(\xi)_{\eta}\|_{\mathbb{M}})$. Since the mapping $y$ is locally Lipschitz continuous, $r_{V}F_{\eta}(\xi_{\phi})$ is of order $o(\|x_{\phi}-\xi_{\phi}\|_{\mathbb{X}})$ as $x_{\phi}\to\xi_{\phi}$.
Thus, we have to discuss the remaining two terms
$-B_{y(x),\eta}(F_{\eta}(x_{\phi}))\delta y_{\eta}-((V_{y(x),\eta}^{-1})^{\prime}(y(x)_{\eta})\delta y_{\eta})F_{\eta}(\xi_{\phi}).$ (14)
Using the bilinearity of $B_{y(x),\eta}$ we can write (14) as
$-B_{y(x),\eta}(F_{\eta}(x_{\phi})-F_{\eta}(\xi_{\phi}))\delta y_{\eta}-\big{[}B_{y(x),\eta}(F(\xi_{\phi}))\delta y_{\eta}+((V_{y(x),\eta}^{-1})^{\prime}(y(x)_{\eta})\delta y_{\eta})(F_{\eta}(\xi_{\phi}))\big{]}$
Since $F$ is Newton-differentiable and $x\mapsto B_{y(x),\eta}$ is locally bounded, the first term is of order $\mathcal{O}(\|x_{\phi}-\xi_{\phi}\|\|\delta y_{\eta}\|)$, and thus, by Lipschitz continuity of $y$, of order $o(\|x_{\phi}-\xi_{\phi}\|)$. If $V^{-1}_{y}$ is consistent with the connection map $Q$, i.e. $B_{y(x),\eta}=-(V_{y(x),\eta}^{-1})^{\prime}(y(x)_{\eta})$, or $F(\xi)=0$ the second term in braces vanishes. Thus, the numerator of $(i)$ is of order $o(\|x_{\phi}-\xi_{\phi}\|_{\mathbb{X}})$ as $x_{\phi}\to\xi_{\phi}$. Finally, using Prop. 5.5 to estimate $d_{\mathcal{X}}(x,\xi)$, we obtain the desired result.
$(ii)$: Let $F$ be strictly differentiable at $a\in\mathcal{X}$. Let $x,\xi\in\mathcal{X}$. If $V^{-1}_{y}$ is consistent with the connection map $Q$ we have
$B_{y(a),\eta}(F_{\eta}(a_{\phi}))\delta y_{\eta}=-((V_{y(a),\eta}^{-1})^{\prime}(y(a)_{\eta})\delta y_{\eta})F_{\eta}(a_{\phi}).$
Otherwise, we assume that $F(a)=0$. In both cases our representation reads (the sign in the second line depends on these cases):
$(Id_{E_{y(a)}}-V_{y(a),\eta}^{-1}(y(x)_{\eta}))(F_{\eta})^{\prime}(a_{\phi})\delta x_{\phi}$
±$\pm((V_{y(a),\eta}^{-1})^{\prime}(y(a)_{\eta})\delta y_{\eta})(F_{\eta}(a_{\phi})-F_{\eta}(\xi_{\phi}))$
$+(F_{\eta})^{\prime}(a_{\phi})r_{R}-V_{y(a),\eta}^{-1}(y(x)_{\eta})r_{F}-F_{\eta}(\xi_{\phi})r_{V}.$
@@
Since $(F_{\eta})^{\prime}(a_{\phi})$ is fixed, we obtain that $(F_{\eta})^{\prime}(a_{\phi})\delta x_{\phi}$ is of order $\mathcal{O}(\|x_{\phi}-\xi_{\phi}\|_{\mathbb{X}})$. Using that $V_{y(a),\eta}^{-1}(y(x)_{\eta})\to Id_{E_{y(a)}}$ for $y(x)\to y(a)$, and $x\mapsto y(x)$ is continuous, we get that $Id_{E_{y(a)}}-V_{y(a),\eta}^{-1}(y(x)_{\eta})$ is of order $\omega(\|x_{\phi}-a_{\phi}\|_{\mathbb{X}})$ as $x\to a$. Thus, the first line vanishes for $x,\xi\to a$.
By strict differentiability, we have that $F_{\eta}(a_{\phi})-F_{\eta}(\xi_{\phi})$ is of order $\omega(\|\xi_{\phi}-a_{\phi}\|_{\mathbb{X}})$ as $\xi\to a$. The mapping $y$ is locally Lipschitz continuous, and thus $\delta y_{\eta}$ is of order $\mathcal{O}(\|x_{\phi}-\xi_{\phi}\|_{\mathbb{X}})$. In addition $(V_{y(a),\eta}^{-1})^{\prime}(y(a)_{\eta})$ is bilinear. Thus, the second line also vanishes for $x,\xi\to a$.
The mappings $R_{a}^{-1}$ and $V^{-1}_{y}$ are continuously differentiable and thus strictly differentiable. Using in addition that $(F_{\eta})^{\prime}(a_{\phi})$ is constant, $V_{y(a),\eta}^{-1}(y(x)_{\eta})$ is locally bounded and $F$ is strictly differentiable, all the terms in the third line vanish for $x,\xi\to a$. By using Prop. 5.5 and Prop. 5.8 this finally yields the desired result $(ii)$. ∎

### v2

@@ 5.1. Local norms and a metric on manifolds
Let $\mathcal{X}$ be a $C^{1}$-Banach manifold, $x_{0}\in\mathcal{X}$ and let $R$ be a $C^{1}$-retraction. Let $(V,\phi)$ be a chart of $\mathcal{X}$ at $x_{0}$ and $\Phi$ be the corresponding natural tangent bundle trivialization.
Then for any $\varepsilon>0$ there exists a neighborhood $U\subset V$ of $x_{0}$, such that the following holds:
  • i)
    A local inverse $R_{z}^{-1}:U\to T_{z}\mathcal{X}$ exists for each $z\in U$. Denote their chart representation by
    $R_{z,\phi}^{-1}:\phi(U)\to\mathbb{X}.$ (12)
  • ii)
    For all $x,y,z\in U$ we have the estimate:
    $\|x_{\phi}-y_{\phi}-(R_{z,\phi}^{-1}(x_{\phi})-R_{z,\phi}^{-1}(y_{\phi}))\|_{\Phi,x_{0}}\leq\varepsilon\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}.$ (13)
Proof.
Application of the inverse function theorem to the chart representation $\widehat{R}_{\phi}$ of the mapping
$\widehat{R}:T\mathcal{X}$ $\to\mathcal{X}\times\mathcal{X}$
$(z,v)$ $\mapsto(z,R_{z}(v))$
yields that $\widehat{R}_{\phi}:\phi(V)\times\mathbb{X}\to\mathbb{X}\times\mathbb{X}$ is a local diffeomorphism, and it holds $(z_{\phi},R_{z,\phi}^{-1}(x_{\phi}))=\widehat{R}_{\phi}^{-1}(z_{\phi},x_{\phi})$. Since $\widehat{R}_{\phi}^{-1}$ is continuously differentiable and thus strictly differentiable we obtain a neighborhood $W\subset\mathbb{X}\times\mathbb{X}$ of $(x_{0,\phi},x_{0,\phi})$, such that for $(z_{\phi},x_{\phi}),(z_{\phi},y_{\phi})\in W$ we have, using $(z_{\phi},x_{\phi})-(z_{\phi},y_{\phi})=(0,x_{\phi}-y_{\phi})$ and denoting by $[v]_{2}$ the second component of $v\in\mathbb{X}\times\mathbb{X}$:
$\|[(\widehat{R}_{\phi}^{-1})^{\prime}(x_{0,\phi},x_{0,\phi})(0,x_{\phi}-y_{\phi})-(\widehat{R}_{\phi}^{-1}(z_{\phi},x_{\phi})-\widehat{R}_{\phi}^{-1}(z_{\phi},y_{\phi}))]_{2}\|_{\Phi,x_{0}}\leq\varepsilon\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}.$ (14)
Taking into account that $(R_{x_{0},\phi}^{-1})^{\prime}(x_{0,\phi})=(R_{x_{0},\phi}^{\prime}(R_{x_{0},\phi}^{-1}(x_{0},\phi)))^{-1}=Id_{\mathbb{X}}$, this in turn implies (13), if we choose $U$, such that for $x,y,z\in U$ we get $(z_{\phi},x_{\phi}),(z_{\phi},y_{\phi})\in W$. ∎
@@
To $x_{0}\in\mathcal{X}$ and $\varepsilon>0$ there exists a neighborhood $U$ of $x_{0}$ such that for all $x,y,z\in U$ the following holds:
@@
Choose a chart $(V,\phi)$ of $\mathcal{X}$ at $x_{0}$ and denote by $\Phi:\pi^{-1}(V)\to\mathbb{X}\times\mathbb{X}$ the natural vector bundle chart on the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$. Let $x_{0,\phi},\ x_{\phi},\ y_{\phi}$ and $z_{\phi}$ be representatives of $x_{0},x,y$ and $z$ in the chart. Let $R_{z,\phi}^{-1}$ be the chart representation of $R_{z}^{-1}$ according to (12). We show that
$\lim_{x_{\phi},y_{\phi},z_{\phi}\rightarrow x_{0,\phi}}\frac{\|R_{z,\phi}^{-1}(x_{\phi})-R_{z,\phi}^{-1}(y_{\phi})\|_{\Phi,x_{0}}}{\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}}=1.$
Using (13) and the inverse triangle inequality, we obtain that for any $\varepsilon>0$ there exists a neighborhood $V$ of $x_{0}$, such that
$\big{|}\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}-\|R_{z,\phi}^{-1}(x_{\phi})-R_{z,\phi}^{-1}(y_{\phi})\|_{\Phi,x_{0}}\big{|}$ $\leq\varepsilon\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}.$
By possibly shrinking $V$, we obtain the desired assertion by Prop. 5.6. ∎
@@ 5.2. Newton- and Strict Differentiability in Geometric Form
In order to analyse local convergence of Newton’s method, we want to prove a result about a Newton- or strictly differentiable mapping $F:\mathcal{X}\to\mathcal{E}$ between a manifold $\mathcal{X}$ and a vector bundle $p:\mathcal{E}\to\mathcal{Y}$, which can be stated independently of the choice of charts.
@@
$y=p\circ F:\mathcal{X}\to\mathcal{Y},\;y(x):=p(F(x))\in\mathcal{Y},$
which computes the base point of $F(x)\in\mathcal{E}$ for any $x\in\mathcal{X}$, is locally Lipschitz continuous. We can use the linear map $V^{-1}_{y}(y_{0})\in L(E_{y_{0}},E_{y})$ given by a vector back-transport $V_{y}^{-1}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ to transport elements $F(x_{0})\in E_{y_{0}}$ into the fibre $E_{y}$. Such a vector back-transport may be consistent to a connection map $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$, i.e., according to Def. 2.4, we may have $Q_{e}=\langle V_{y}^{-1}\rangle^{\prime}(e),\;e\in\mathcal{E}$.
@@
Let $F:\mathcal{X}\to\mathcal{E}$. Let $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ be a vector back-transport, $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$ be a connection map and $R:T\mathcal{X}\to\mathcal{X}$ be a $C^{1}$-retraction.
@@
    Let $F:\mathcal{X}\to\mathcal{E}$ be Newton-differentiable at $x_{0}\in\mathcal{X}$ with respect to $F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{E}))$. If $V^{-1}_{y}$ is consistent with $Q$ or $F(x_{0})=0$, then it holds
    $\underset{x\to x_{0}}{\lim}\frac{\|(Q_{F(x)}\circ F^{\prime}(x))(R_{x}^{-1}(x)-R_{x}^{-1}(x_{0}))-(F(x)-V_{y(x)}^{-1}(y(x_{0}))F(x_{0}))\|_{E_{y(x)}}}{d_{\mathcal{X}}(x,x_{0})}=0.$
@@
    Let $F$ be strictly differentiable at $x_{0}\in\mathcal{X}$. If $V^{-1}_{y}$ is consistent with $Q$ or $F(x_{0})=0$ then for every $\varepsilon>0$ there exists a neighborhood $W$ of $x_{0}$ such that for all $x,\xi\in W$ it holds
    $\|(Q_{F(x_{0})}\circ F^{\prime}(x_{0}))(R_{x_{0}}^{-1}(x)$ $-R_{x_{0}}^{-1}(\xi))-(V^{-1}_{y(x_{0})}(y(x))F(x)-V_{y(x_{0})}^{-1}(y(\xi))F(\xi))\|_{E_{y(x_{0})}}$
    $<\varepsilon\|R_{x_{0}}^{-1}(x)-R_{x_{0}}^{-1}(\xi)\|_{x_{0}}.$
@@
Choose a chart $(U,\phi)$ of $\mathcal{X}$ and a vector bundle chart $\eta$ of $\mathcal{E}$. Let $u,v,w\in U$. In the following, we use $\phi$ or $\eta$ as an index to denote the chart representation of elements of $\mathcal{X}$ or $\mathcal{E}$, respectively. We begin by discussing the numerators in $(i)$ and $(ii)$ in a unified way and collecting all remainder terms which occur. Then, for the proof of $(i)$ we set $u=v=x$ and fix $w=x_{0}$. For the proof of $(ii)$ we will fix $u=x_{0}$ and consider $v=x,w=\xi$. With these notations the remainder terms read in both cases
$r(u,v,w)=(Q_{F(u)}\circ F^{\prime}(u))(R_{u}^{-1}(v)-R_{u}^{-1}(w))-\big{(}V^{-1}_{y(u)}(y(v))F(v)-V_{y(u)}^{-1}(y(w))F(w)\big{)}.$
Let $R_{u,\phi}^{-1}$ be a representation of the inverse retraction in charts given by (12). Since $R_{u}^{-1}$ is a $C^{1}$-mapping with $(R_{u}^{-1})^{\prime}(u)=R_{u}^{\prime}(R_{u}^{-1}(u))^{-1}=Id_{T_{u}\mathcal{X}}$, we obtain
$R_{u,\phi}^{-1}(v_{\phi})-R_{u,\phi}^{-1}(w_{\phi})=(R^{-1}_{u,\phi})^{\prime}(u_{\phi})(v_{\phi}-w_{\phi})+r_{R^{-1}}=(v_{\phi}-w_{\phi})+r_{R^{-1}}$
where $r_{R^{-1}}$ is a remainder term. Setting $\delta y_{\eta}:=y(v)_{\eta}-y(w)_{\eta}\in\mathbb{Y}$, we obtain the chart representation of
$(Q_{F(u)}\circ F^{\prime}(u))(R_{u}^{-1}(v)-R_{u}^{-1}(w))$
@@
$(F_{\eta,\phi})^{\prime}(u_{\phi})(v_{\phi}-w_{\phi}+r_{R^{-1}})-B_{y(u),\eta}(F_{\eta,\phi}(u_{\phi}))\delta y_{\eta}.$ (15)
Next, we consider the representation of $V_{y(u)}^{-1}(y(v))F(v)-V_{y(u)}^{-1}(y(w))F(w)$, which is of the form
$V^{-1}_{y(u),\eta}(y(v)_{\eta})F_{\eta,\phi}(v_{\phi})-V^{-1}_{y(u),\eta}(y(w)_{\eta})F_{\eta,\phi}(w_{\phi}).$ (16)
Using the differentiability of the mapping $V_{y,\eta}^{-1}:\mathbb{Y}\to L(\mathbb{E},\mathbb{E})$, we can write:
$V_{y(u),\eta}^{-1}(y(v)_{\eta})-V_{y(u),\eta}^{-1}(y(w)_{\eta})=(V_{y(u),\eta}^{-1})^{\prime}(y(u)_{\eta})\delta y_{\eta}+r_{V}$
where $r_{V}$ denotes a remainder term. Inserting this into (16) we obtain by linearity of $V_{y(u)}^{-1}(y(v))$
[…]
Writing the difference $F_{\eta,\phi}(v_{\phi})-F_{\eta,\phi}(w_{\phi})$ as
$F_{\eta,\phi}(v_{\phi})-F_{\eta,\phi}(w_{\phi})=(F_{\eta,\phi})^{\prime}(u_{\phi})(v_{\phi}-w_{\phi})+r_{F}$
@@
$r(u,v,w)\sim r_{\eta,\phi}(u_{\phi},v_{\phi},w_{\phi})$ $=(Id_{E_{y(u)}}-V_{y(u),\eta}^{-1}(y(v)_{\eta}))(F_{\eta,\phi})^{\prime}(u_{\phi})(v_{\phi}-w_{\phi})$ (18)
$-B_{y(u),\eta}(F_{\eta,\phi}(u_{\phi}))\delta y_{\eta}-((V_{y(u),\eta}^{-1})^{\prime}(y(u)_{\eta})\delta y_{\eta})F_{\eta,\phi}(w_{\phi})$
$+(F_{\eta,\phi})^{\prime}(u_{\phi})r_{R^{-1}}-V_{y(u),\eta}^{-1}(y(v)_{\eta})r_{F}-r_{V}F_{\eta,\phi}(w_{\phi}).$
We will discuss the three lines of (18) separately.
@@
In the Newton-differentiable case, where $u=v=x$ we have $V_{y(u)}^{-1}(y(v))=Id_{E_{y}}$ and thus, the first line vanishes. In the strictly differentiable case $(F_{\eta,\phi})^{\prime}(u_{\phi})$ is fixed. Thus, we obtain that $(F_{\eta,\phi})^{\prime}(u_{\phi})(v_{\phi}-w_{\phi})$ is of order $\mathcal{O}(\|v_{\phi}-w_{\phi}\|_{\mathbb{X}})$. Using that $V_{y(u),\eta}^{-1}(y(v)_{\eta})\to Id_{E_{y(u)}}$ for $y(v)\to y(u)$, and $x\mapsto y(x)$ is continuous, we get that $Id_{E_{y(u)}}-V_{y(u),\eta}^{-1}(y(v)_{\eta})$ tends to 00 in the operator norm as $v\to u$. Thus, in both cases the first line of (18) vanishes in the desired order for all choices of $u,v,w$ corresponding to the respective differentiability concepts.
Next, we have to discuss the second line of (18) using our consistency assumption or $F(x_{0})=0$. If $V^{-1}_{y}$ is consistent with the connection map $Q$ we have
$B_{y(u),\eta}(F_{\eta,\phi}(u_{\phi}))\delta y_{\eta}=-((V_{y(u),\eta}^{-1})^{\prime}(y(u)_{\eta})\delta y_{\eta})F_{\eta,\phi}(u_{\phi}).$
Thus, in this case the second line of (18) simplifies to
$((V_{y(u),\eta}^{-1})^{\prime}(y(u)_{\eta})\delta y_{\eta})(F_{\eta,\phi}(u_{\phi})-F_{\eta,\phi}(w_{\phi})).$ (19)
Alternatively, we assume that $F(x_{0})=0$. In the Newton differentiable case this yields $F_{\eta,\phi}(w_{\phi})=0$, in the strictly differentiable case we have that $F_{\eta,\phi}(u_{\phi})=0$. Thus, we can rewrite the second line of (18) in the Newton-differentiable case as
$-B_{y(u),\eta}(F_{\eta,\phi}(u_{\phi})-F_{\eta,\phi}(w_{\phi}))\delta y_{\eta},$ (20)
and in the strictly differentiable case as
$((V_{y(u),\eta}^{-1})^{\prime}(y(u)_{\eta})\delta y_{\eta})(F_{\eta,\phi}(u_{\phi})-F_{\eta,\phi}(w_{\phi})).$ (21)
In all the above cases (19)-(21), using Lipschitz continuity of $F_{\eta,\phi}$, bilinearity of $(V_{y(u),\eta}^{-1})^{\prime}(y(u)_{\eta})$ or of $B_{y(u),\eta}$, and Lipschitz continuity of the mapping $x\mapsto y(x)$, implying $\delta y_{\eta}=\mathcal{O}(\|v_{\phi}-w_{\phi}\|_{\mathbb{X}})$, the second line also vanishes in the desired order. Finally, we discuss the third line of (18), where the remainder terms $r_{R^{-1}}$, $r_{V}$, and $r_{F}$ vanish in both cases in the desired order, since the mappings $R_{u}^{-1}$ and $V^{-1}_{y}$ are continuously differentiable and $F$ is Newton- respectively strictly differentiable. In addition $(F_{\eta,\phi})^{\prime}(u_{\phi})$ is locally bounded by assumption in the Newton-differentiable case and fixed in the strict differentiable case, $V_{y(u),\eta}^{-1}(y(v)_{\eta})$ is locally bounded and $F_{\eta,\phi}(w_{\phi})$ is locally bounded since $w$ is fixed in the Newton-differentiable case and by continuity in the strictly differentiable case. Thus, the third line of (18) also vanishes in the desired order if we choose $u,v,w$ corresponding to the respective differentiability concepts.
In summary, we get the following smallness result. For each $\varepsilon>0$ there exists a neighbourhood $U_{\varepsilon}$ of $x_{0}$, such that
$\|r_{\eta,\phi}(u_{\phi},v_{\phi},w_{\phi})\|_{\mathbb{E}}\leq\varepsilon\|v_{\phi}-w_{\phi}\|_{\mathbb{X}}$
for all choices of $u,v,w\in U_{\varepsilon}$ corresponding to the respective differentiability concept. By the equivalences of norms, established in Prop. 5.6 and Prop. 5.9 in addition, we obtain the desired results. ∎